the length of a rectangle is 6cm longer than th width. The perimeter of the rectangle is 92 cm. What is the length and the width of the rectangle?
Answer by ikleyn(53763)

.
x + (x+6) = 92/2


2x + 6 = 46


2x = 46-6 = 40


x = 40/2 = 20.


Answer.  The width is 20 cm;  the length is 26 cm

Solved.